### v2

@@ 2.4 Pathological Image Retrieval
Image retrieval techniques could match the new patient pathology images to a curated database of previously diagnosed cases, providing pathologists with a novel tool to enhance diagnostic accuracy. Through visual inspection and comparison of similar historical cases, pathologists can leverage image search functionality to enhance their diagnostic decision-making. This approach allows pathologists to go beyond the capabilities of individual AI algorithms and instead harness the collective knowledge distilled from large repositories of validated pathology images and expert diagnoses [57, 58, 59]. In this study, we employ the CRC-100K dataset [60] for conducting pathological image retrieval tasks. This dataset comprises 100,000 non-overlapping 224224 patches extracted from 86 human cancer tissue slides stained with H&E for training purposes. Additionally, it includes 7,180 images with 224224 pixels extracted from 50 patients diagnosed with colorectal adenocarcinoma for testing. The dataset consists of multiple classes, including Adipose (ADI, 11,745 ROIs), Background (BACK, 11,413 ROIs), Debris (DEB, 11,851 ROIs), Lymphocytes (LYM, 12,191 ROIs), Mucus (MUC, 9,931 ROIs), Smooth muscle (MUS, 14,128 ROIs), Normal colon mucosa (NORM, 9,504 ROIs), Cancer-associated stroma (STR, 10,867 ROIs), and Colorectal adenocarcinoma epithelium (TUM, 15,550 ROIs). For training and evaluation, we utilize the official train-test split, with 100,000 samples for training and 7,180 samples for testing.
The experimental results for ROI retrieval are depicted in Figure 6, and the detailed results can be found in Extended Data Table 33, showcasing the Top-1, Top-3, and Top-5 accuracy achieved by different foundation models. We observe that the GPFM model achieved the second-best Top-1 accuracy with a value of 0.906 (-0.5%, UNI). However, GPFM outperforms other models in terms of Top-3 and Top-5 accuracy, achieving values of 0.993 (+1.2%, UNI) and 0.995 (+1.2%, UNI) respectively. Furthermore, to gain insights into the clustering effect and the quality of feature representation, we employ t-SNE (t-Distributed Stochastic Neighbor Embedding) [61] to reduce the dimensionality of features extracted by various foundation models into 2D embedding. This visualization technique allows us to observe the clustering patterns of different classes within the dataset. As illustrated in Extended Data Figure 12, the features extracted by the GPFM model are clustered more tightly and the query image is also located within the candidate cluster, indicating a better clustering effect. This observation suggests that the GPFM model has superior feature representation capabilities in capturing the intrinsic patterns and structures present in the data.
@@ 2.5 Pathological Images VQA
Visual Question Answering (VQA) is an exciting field of artificial intelligence that aims to enable machines to answer questions about visual content. In the domain of pathology, VQA systems can be particularly powerful, allowing clinicians and researchers to quickly and accurately extract relevant information from medical images. To evaluate the effectiveness of foundation models in pathology VQA, we utilize the PathVQA dataset [62], which is the largest and most widely used dataset in the pathology domain for VQA tasks. The dataset consists of 32,799 image-question-answer triplets, divided into three subsets: a training set (50%) containing 16,400 triplets used for model training, a validation set (30%) comprising 9,840 triplets for hyperparameter tuning and overfitting prevention, and a test set (20%) including 6,560 triplets for final model performance evaluation. The performance of different foundation models on open-ended and close-ended VQA problems is presented in Figure 6, and detailed results can be found in Extended Data Table 34. Our model achieved the second-best performance, performing only slightly lower than CONCH. It is worth noting that CONCH is a vision-language foundation model trained on million-scale image-text pairs, which naturally gives it an advantage in VQA tasks. The results demonstrate the significant potential of our approach for VQA tasks compared with other pure vision foundation models. Additionally, we have provided visualizations of the query image, questions, and answers generated by different foundation models in Figure 6.c and e. As shown in the figure, the GPFM and CONCH foundation models consistently produced more reliable answers compared to the other models evaluated.
Through unified knowledge distillation, the knowledge that CONCH acquired from the million image-text pairs can be distilled into GPFM without using any of the original image-text pair data. The performance of GPFM demonstrates the potential of leveraging textual knowledge without directly utilizing text data.
@@ 2.6 Pathology Report Generation
Pathology reports are essential components of the healthcare system, providing critical information to clinicians and patients about the diagnosis, prognosis, and treatment of various medical conditions. These reports summarize the findings from pathological examinations, such as biopsies, cytology samples, and surgical specimens, and play a vital role in guiding clinical decision-making. Traditionally, pathology reports are written manually by pathologists and their teams, a time-consuming and labor-intensive process. Recent advancements in natural language processing (NLP) and machine learning have enabled the development of automated pathology report generation systems, which can dramatically improve the efficiency and consistency of this critical task [63, 64, 65]. To evaluate the effectiveness of foundation models in pathology report generation, we utilize the TCGA WSI-report dataset curated by Guo et al. [63]. This dataset consists of 7,690 pairs of WSIs and their corresponding diagnosis reports, covering various diseases from different primary sites. In the task of WSI report generation, foundation models are employed to provide representations of WSIs as inputs to the report generation model. We use HistGen [63] as the baseline model, which takes the WSI features extracted by foundation models as input and generates the corresponding diagnosis report. For training and evaluation, we split the dataset into train-validation-test cohorts with a ratio of 7:1:2 (5,383: 769: 1,538 slides). The performance of each foundation model in report generation is presented in Figure 7 and the detailed results are reported in Extended Data Table 35.
The experimental results demonstrate that Phikon achieved the best performance across all six metrics, while GPFM achieved comparable performance and ranked as the second-best model. This is quite surprising that vision foundation models (e.g., Phikon and GPFM) achieved much better performance in this task compared to vision-language foundation models such as CONCH and PLIP. A potential reason behind is that PLIP and CONCH were only trained with short descriptions or captions of pathological images, without access to the global contextual information. This may have made the text-image pairs less effective for this task compared to the VQA tasks they were originally designed for. It is worth noting that the Phikon model was pretrained exclusively on data from TCGA. While this may enhance Phikon’s efficacy on tasks specifically utilizing data derived from the TCGA, it also limits Phikon’s generalization ability on tasks outside of the TCGA domain. To leverage the complementary strengths of existing models, the proposed unified knowledge distillation approach can distill the capabilities of Phikon in report generation into the GPFM. This synergistic integration allows us to combine the respective strengths of these foundation models, leading to the development of a more generalizable model.
@@ 2.7 The Effectiveness of Expert Knowledge Distillation
In the self-supervised learning framework proposed in this study, we introduced a unified knowledge distillation model to facilitate the transfer of knowledge from off-the-shelf foundation models to GPFM during the pretraining stage. To assess the effectiveness of this module, we conducted an experiment where we removed the Expert Knowledge Distillation module, resulting in a modified self-supervised learning framework known as DINOv2 [56]. We trained both DINOv2 and GPFM on the same dataset and evaluated their performance in tissue classification tasks. The experimental results are presented in Figure 8. More details can be found in Extended Data Table 36. The experimental results clearly demonstrate the positive impact of Expert Knowledge Distillation on the performance of the models across 12 tasks. Not only did the individual task performances improve significantly, but the average performance also exhibited enhancement, with notable improvements in all three metrics. The AUC increased by 0.6%, the weighted F1 score improved by 1.8%, and the balanced accuracy showed an increase of 1.8%. These findings provide strong evidence for the effectiveness of transferring knowledge from off-the-shelf pathology foundation models through the proposed knowledge distillation learning framework.
@@ 3 Discussion
In this study, we construct the most comprehensive benchmark for CPath tasks to date, to the best of our knowledge. Additionally, we introduce GPFM, a generalizable foundation model designed for a broad spectrum of CPath tasks. To enhance the model’s versatility, we propose a unified knowledge distillation pretraining framework, which effectively consolidates expertise from a variety of existing models. This innovative approach ensures that GPFM can adapt and excel across different CPath tasks. To further maximize the diversity of data used for pretraining, we gathered 190 million images sourced from 47 sources, spanning 34 major tissue types. This rich dataset, combined with our advanced pretraining methodology, empowers GPFM to surpass current foundation models in performance across six major categories (WSI classification, survival analysis, ROI classification, image retrieval, VQA, and report generation) of CPath tasks, comprising a total of 39 specific tasks. Unlike other models that demonstrate proficiency in narrow domains—such as UNI [33], which specializes in ROI classification, and Phikon [32], which excels in survival analysis—GPFM showcases exceptional generalization, outperforming its counterparts across a wide array of CPath challenges by combining the strengthens of expert models.
@@
Recently, several vision-language [35, 36] and pure vision [33, 32, 37, 66] pathology foundation models have been developed. However, the overall performance of these existing foundation models is unclear due to the absence of a comprehensive benchmark. Our analysis reveals that no single existing model consistently exhibits the best performance. This is likely because each foundation model is trained using distinct datasets and pretraining strategies, leading to model-specific advantages for particular domains and datasets. The root of a model’s generalization ability lies in the diversity of the training data. Unfortunately, gathering extremely large-scale diverse datasets, especially for sensitive medical data, is very difficult due to security and privacy concerns. Therefore, it is almost impossible to access and utilize all the data used to develop the existing foundation model. Although accessing the original private training data is limited, the pretrained models themselves are available. Since the knowledge of the pretrained models is derived from the training data, we can indirectly leverage this knowledge by using a unified knowledge distillation framework. It provides a feasible method to integrate knowledge from a large number of existing models under the premise of limited data and protecting data privacy, which has better feasibility and scalability in clinical practice. The significantly greater generalization ability of GPFM compared to existing foundation models, suggests that transferring knowledge from one existing model to another may be a more viable path to further advancing pathology foundation models in the future, especially given the challenges of assembling large-scale diverse medical datasets.
@@
This study also has some limitations. While we established a most comprehensive benchmark covering 39 tasks and demonstrated that the GPFM was the most generalizable foundation model, the external validation was limited to only seven tasks. This set of external validation tasks could be expanded in future work to further assess the generalizability of the models. It is worth highlighting that, across the seven external validation tasks, the GPFM achieved an average rank of 1.57, while the second-best performing model, UNI, achieved an average rank of 2.79. This result is consistent with the conclusions drawn from the experiments across the entire benchmark, highlighting the superior generalizability of GPFM compared to other models. Nevertheless, we also recognize that current off-the-shelf foundation models still exhibit potential in specific tasks, such as Phikon for report generation using TCGA data. This illustrates that the proposed unified knowledge distillation approach is not perfect and has room for improvement. Future research should concentrate on developing sophisticated methodologies to effectively distill and incorporate expert knowledge into one model, maximizing their potential across a broader spectrum of tasks. For example, further expanding the model’s parameter size to enhance its adaptability, facilitating a more comprehensive assimilation of knowledge from diverse foundation models. Additionally, the current GPFM is an unimodal foundation model, which limits its ability to effectively handle cross-modal tasks such as VQA. Given the prevalence of multimodal data in pathology, encompassing WSIs, reports, and genomic data, the development of a multi-modal pathology foundation model is more attractive. Such a model would be more adept at integrating heterogeneous information, offering a more holistic understanding of patient data and enhancing diagnostic accuracy. Last, the GPFM and existing foundation models function as ROI-level feature extractors. However, the majority of CPath tasks demand analysis at the whole-slide level. Recent advancements in long sequence modeling techniques [67, 68], exemplified by MambaMIL [69], have demonstrated success in WSI classification tasks. Developing whole slide-level foundation model could avoid the reliance on MIL frameworks, potentially leading to significant strides in CPath by providing more contextually rich interpretations of tissues.
@@ 4 Methods
4.1 Foundation Model Pretraining
CPath has emerged as a groundbreaking field that synergizes the power of AI with the expertise of pathologists, revolutionizing the practice of diagnosing and analyzing diseases. At the core of this transformative discipline lies the foundation model, which serves as the backbone for a wide range of applications in pathology.
While there exist some readily available foundation models such as Ctranspath (pretrained on 32K TCGA slides) [37] and UNI [33] (pretrained on 100K private slides), the utilization of public data remains incomplete, and the evaluation of these models in CPath tasks is inadequate.
The limited diversity of primary sites in the pretraining slides also restricts the adaptability of current foundation models for public CPath benchmarks.
To facilitate the advancement of CPath, we meticulously curated a comprehensive dataset comprising 47 publicly available histopathology datasets, encompassing a wide spectrum of 34 distinct tissue types for pretraining and downstream task evaluation.
The primary sites of tissues are reported in Extended Data Table 41.
Leveraging this large-scale dataset, we developed a self-supervised learning approach with unified knowledge distillation to construct a foundation model that surpasses existing models.
Dataset Preparation. To boost the performance of foundation models, diverse datasets with various tissues are necessary.
@@ 4.1 Foundation Model Pretraining
It is worth noting that we did not scale the patches to a uniform resolution, opting instead to use the original resolution of each WSI. This approach was implemented to increase the robustness of the foundation models to varying resolutions.
@@
The dataset used for pretraining consists of 72,280 WSIs and a total of 190,212,668 patches.
Pretraining with Self and Expert Knowledge Distillation. In CPath, current foundation models typically rely on state-of-the-art self-supervised pretraining (SSL) methods, such as DINOv2 [56] and iBOT [72]. These methods are applied directly to either private or public datasets.
@@
Due to its larger training dataset and more powerful SSL methods, UNI outperforms Phikon in various tasks. However, UNI still lags behind other foundation models in tasks related to text analysis and survival analysis due to its pretraining strategy and limited coverage of primary sites.
To address the limitations of current foundation models and further enhance their performance, we propose a novel pretraining strategy involving Unified Knowledge Distillation.
The framework of our proposed pretraining method is illustrated in Figure 1.c.
Similar to DINOv2, we employ teacher-student networks with masking image modeling (MIM) loss [73] and DINO (self-distillation) [74, 72] loss to optimize the student network. Specifically, given an input image , we obtain two augmented views, and .
@@
However, this strategy fails to leverage the knowledge from existing vision foundation models, such as UNI and vision-language foundation models like CONCH [35], which restricts their applicability to different tissue types. To transfer the knowledge from established pathology foundation models, we introduce an Expert Knowledge Distillation module to distill knowledge into the student network [38, 75].
Specifically, we adopt three foundation models that excel in classification tasks (UNI), survival analysis tasks (Phikon), and visual question answering tasks (CONCH).
To achieve distillation, we use the student network to encode the global views and and extract the [CLS] and [PATCH] tokens. Additionally, we employ the off-the-shelf foundation models (UNI, Phikon, CONCH) to obtain their respective [CLS] and [PATCH] tokens. For aligning the class tokens, we utilize cosine similarity. As for the patch token alignment, we employ both cosine similarity and smooth L1 distance. The pseudo-code for this process is outlined in Algorithm 1.
@@
Baselines. To evaluate the performance of our foundation model, GPFM, we conducted a comprehensive evaluation by comparing it with other vision foundation models, namely Ctranspath [37], Phikon [32], and UNI [33], as well as visual-language foundation models PLIP [36] and CONCH [35].
As a baseline, we also compared these foundation models with a ResNet50 [76] pretrained on the ImageNet dataset [77].
@@
It is important to note that, for all images used in the downstream tasks, the feature extraction operations were performed on images resized to 224224, unless specified otherwise.

### v3

@@ 2.4 Pathological Image Retrieval
Image retrieval techniques could match the new patient pathology images to a curated database of previously diagnosed cases, providing pathologists with a novel tool to enhance diagnostic accuracy. Through visual inspection and comparison of similar historical cases, pathologists can leverage image search functionality to enhance their diagnostic decision-making. In this study, we employ the CRC-100K dataset [49] for conducting pathological image retrieval tasks.
The experimental results (Fig. 6a, Extended Data Table 37) show that the GPFM model achieved the second-best Top-1 accuracy with a value of 0.906 (-1.9%, Prov-Gigapath). However, GPFM outperforms other models in terms of Top-3 and Top-5 accuracy, achieving values of 0.993 (+0.5%, Prov-Gigapath) and 0.995 (+0.2%, Prov-Gigapath), respectively. To further explore the clustering effect and feature representation ability, we utilized t-Distributed Stochastic Neighbor Embedding (t-SNE) [50] to project the features extracted by GPFM into a 2D embedding space. The categories are well clustered, further illustrating that the features are highly discriminative (Fig. 6b). We also visualized the feature distribution of other FMs (Extended Data Fig. 10). The features extracted by the GPFM are clustered more tightly and the query image is also located within the candidate cluster, indicating a better clustering effect. This observation suggests that the GPFM has superior feature representation capabilities in capturing the intrinsic patterns and structures present in the data.
@@ 2.5 Pathological Images VQA
VQA is an exciting field of artificial intelligence that aims to enable machines to answer questions about visual content. In the domain of pathology, VQA systems can be particularly powerful, allowing clinicians and researchers to quickly and accurately extract relevant information from medical images.
For the patch-level VQA task, our model achieved the second-best performance, with results only slightly lower than those of CONCH (Fig. 6c, Extended Data Table 38). It is important to note that CONCH is a vision-language FM trained on millions of image-text pairs, which inherently provides it with an advantage in VQA tasks. Despite this, our results highlight the substantial potential of our approach compared to other pure vision FMs. To further illustrate the capabilities of our model, we visualized the query images, questions, and answers generated by different FMs (Fig. 6d and 6e). As demonstrated in the figures, both GPFM and CONCH consistently produced more reliable and accurate answers compared to the other models.
Moreover, in the WSI-level VQA task [51], our model achieved the best or second-best performance across 6 out of 7 metrics, demonstrating performance comparable to the slide-level FM CHIEF (Extended Data Fig. 12 and Table 39). These results, combined with the patch-level findings, underscore the effectiveness of unified knowledge distillation. Specifically, the knowledge acquired by CONCH from millions of image-text pairs can be successfully distilled into GPFM without requiring access to the original image-text pair data. The strong performance of GPFM highlights the potential of leveraging textual knowledge indirectly, without the need for direct utilization of text data, thereby offering a promising direction for future research in VQA tasks.
@@ 2.6 Pathology Report Generation
Pathology reports are essential components of the healthcare system, providing critical information to clinicians and patients about the diagnosis, prognosis, and treatment of various medical conditions. These reports summarize the findings from pathological examinations, such as biopsies, cytology samples, and surgical specimens, and play a vital role in guiding clinical decision-making. Traditionally, pathology reports are written manually by pathologists and their teams, a time-consuming and labor-intensive process. Recent advancements in natural language processing (NLP) and machine learning have enabled the development of automated pathology report generation systems, which can dramatically improve the efficiency and consistency of this critical task [52, 53, 54]. To assess the effectiveness of FMs in this domain, we evaluated their performance on the TCGA WSI-report dataset, curated by Guo et al. [52], and the PatchGastricADC22 [55] dataset.
The experimental results demonstrate that Phikon achieved the best performance across all six metrics, while GPFM achieved comparable performance and ranked as the second-best model on both tasks (Fig. 6f, Extended Data Table 40 and 42). It is quite surprising to observe that vision FMs (e.g., Phikon and GPFM) performed much better in this task than vision-language FMs such as CONCH and PLIP. This performance gap may be attributed to PLIP and CONCH’s training paradigm, which relied solely on short descriptions or captions of pathological images without access to global contextual information. Consequently, these text-image pairs proved less effective for comprehensive report generation compared to their original VQA task applications. The examples of generated reports shown in Extended Data Fig. 13 and 14 certify this assumption.
To further validate these findings, we conducted stratified report generation analyses by stratifying the TCGA WSI-report dataset by major cancer types, i.e., breast, lung, and kidney cancers, for independent evaluation. Results (Extended Data Table 41 and Fig. 11a-c) reveal that Phikon keeps its superiority in breast and lung cancer report generation, yet is slightly outperformed by our GPFM in kidney cancer report generation. To leverage the complementary strengths of existing FMs, the proposed unified knowledge distillation approach can distill the capabilities of Phikon in report generation into the GPFM. This synergistic integration allows us to combine the respective strengths of these FMs, leading to the development of a more generalizable model. To further assess clinical relevance, an experienced pathologist evaluated the diagnostic reports using a four-tier scoring system (Extended Data Fig. 11d). The blinded human-based evaluation results demonstrate GPFM’s superior performance, achieving the highest average scores across breast, lung, and kidney cancer reports (Extended Data Table 43 and Fig. 11a-c). These expert-validated results underscore the potential of our unified knowledge distillation approach to generate clinically meaningful reports that align with pathologists’ diagnostic standards, marking a significant step toward the practical application of AI in pathology workflow automation.
@@ 2.7 The Effectiveness of Expert Knowledge Distillation
In the self-supervised learning framework proposed in this study, we introduced a unified knowledge distillation model to facilitate the transfer of knowledge from off-the-shelf FMs to GPFM during the pretraining stage. To assess the effectiveness of this module, we conducted an experiment where we removed the Expert Knowledge Distillation module, resulting in a modified self-supervised learning framework known as DINOv2 [48]. We trained both DINOv2 and GPFM on the same dataset and evaluated their performance in tissue classification tasks. The experimental results clearly demonstrate the positive impact of Expert Knowledge Distillation on the performance of the models across 12 tasks (Extended Data Fig. 15 and Table 44). The experimental results demonstrated significant improvements not only in the performance of individual tasks but also in the overall average performance, with substantial enhancements observed across all three evaluation metrics. The AUC increased by 0.6%, the weighted F1 score improved by 1.8%, and the balanced accuracy showed an increase of 1.8%. These findings provide strong evidence for the effectiveness of transferring knowledge from off-the-shelf pathology FMs through the proposed knowledge distillation learning framework. However, even with the distillation, GPFM still can not beat vanilla DINOv2 in all tasks such as Chaoyang and BreakHis, illustrating that there is still room for improving the distillation strategy.
@@ 3 Discussion
In this study, we construct the most comprehensive benchmark for CPath tasks to date, to the best of our knowledge. Additionally, we introduce GPFM, a generalizable FM designed for a broad spectrum of CPath tasks. To enhance the model’s versatility, we propose a unified knowledge distillation pretraining framework, which effectively consolidates expertise from a variety of existing models. This innovative approach ensures that GPFM can adapt and excel across different CPath tasks. To further maximize the diversity of data used for pretraining, we gathered 190 million images sourced from 56 sources, spanning 34 major tissue types. This rich dataset, combined with our advanced pretraining methodology, empowers GPFM to surpass current FMs in performance across 72 CPath tasks. Unlike other models that demonstrate proficiency in narrow domains—such as UNI [33], which specializes in WSI classification, and Phikon [32], which excels in report generation—GPFM showcases exceptional generalization, outperforming its counterparts across a wide array of CPath challenges by combining the strengthens of expert models.
@@
Recently, several vision-language [35, 36] and pure vision [33, 32, 37, 56] pathology FMs have been developed. However, the overall performance of these existing FMs is unclear due to the absence of a comprehensive benchmark. Our analysis reveals that no single existing model consistently exhibits the best performance. This is likely because each FM is trained using distinct datasets and pretraining strategies, leading to model-specific advantages for particular domains and datasets. The root of a model’s generalization ability lies in the diversity of the training data. Unfortunately, gathering extremely large-scale diverse datasets, especially for sensitive medical data, is very difficult due to security and privacy concerns. Therefore, it is almost impossible to access and utilize all the data used to develop the existing FM. Although accessing the original private training data is limited, the pretrained models themselves are available. Since the knowledge of the pretrained models is derived from the training data, we can indirectly leverage this knowledge by using a unified knowledge distillation framework. It provides a feasible method to integrate knowledge from a large number of existing models under the premise of limited data and protecting data privacy, which has better feasibility and scalability in clinical practice. The significantly greater generalization ability of GPFM compared to existing FMs, suggests that transferring knowledge from one existing model to another may be a more viable path to further advancing pathology FMs in the future, especially given the challenges of assembling large-scale diverse medical datasets.
@@
This study also has some limitations. We recognize that current off-the-shelf FMs still exhibit potential in specific tasks, such as Phikon for report generation using TCGA data. This illustrates that the proposed unified knowledge distillation approach is not perfect and has room for improvement. Future research should concentrate on developing sophisticated methodologies to effectively distill and incorporate expert knowledge into one model, maximizing their potential across a broader spectrum of tasks. For example, further expanding the model’s parameter size to enhance its adaptability, facilitating a more comprehensive assimilation of knowledge from diverse FMs. Additionally, the current GPFM is an unimodal FM, which limits its ability to effectively handle cross-modal tasks such as VQA. Given the prevalence of multimodal data in pathology, encompassing WSIs, reports, and genomic data, the development of a multi-modal pathology FM is more attractive. Such a model would be more adept at integrating heterogeneous information, offering a more holistic understanding of patient data and enhancing diagnostic accuracy.
@@ 4 Methods
4.1 FM Pretraining
CPath has emerged as a groundbreaking field that synergizes the power of AI with the expertise of pathologists, revolutionizing the practice of diagnosing and analyzing diseases. At the core of this transformative discipline lies the FM, which serves as the backbone for a wide range of applications in pathology.
While there exist some readily available FMs such as Ctranspath (pretrained on 32K TCGA slides) [37] and UNI [33] (pretrained on 100K private slides), the utilization of public data remains incomplete, and the evaluation of these models in CPath tasks is inadequate.
The limited diversity of primary sites in the pretraining slides also restricts the adaptability of current FMs for public CPath benchmarks.
To facilitate the advancement of CPath, we meticulously curated a comprehensive dataset comprising 56 histopathology datasets, encompassing a wide spectrum of 34 distinct tissue types for pretraining and downstream task evaluation (Extended Data Table 50).
Leveraging this large-scale dataset, we developed a self-supervised learning approach with unified knowledge distillation to construct a FM that surpasses existing models.
Dataset Preparation. To boost the performance of FMs, diverse datasets with various tissues are necessary.
@@ 4.1 FM Pretraining
It is worth noting that we did not scale the patches to a uniform resolution, opting instead to use the original resolution of each WSI. This approach was implemented to increase the robustness of the FMs to varying resolutions.
@@
The pretraining data consists of 72,280 WSIs and a total of 190,212,668 patches.
Pretraining with Self and Expert Knowledge Distillation. In CPath, current FMs typically rely on state-of-the-art self-supervised pretraining (SSL) methods, such as DINOv2 [48] and iBOT [59]. These methods are applied directly to either private or public datasets.
@@
Due to larger training dataset and more powerful SSL methods, UNI outperforms Phikon in various tasks. However, UNI still lags behind other FMs in tasks related to text analysis and survival analysis due to its pretraining strategy and limited coverage of primary sites.
To address the limitations of current FMs and further enhance their performance, we propose a novel pretraining strategy involving Unified Knowledge Distillation.
The framework of the proposed pretraining method is similar to DINOv2, we employ teacher-student networks with masking image modeling (MIM) loss [60] and DINO (self-distillation) [61, 59] loss to optimize the student network (Fig. 1c).
Specifically, given an input image , we obtain two augmented views, and .
@@
However, this strategy fails to leverage the knowledge from existing vision FMs, such as UNI and vision-language FMs like CONCH [35], which restricts their applicability to different tissue types.
To facilitate the transfer of knowledge from established pathology FMs, we propose an Expert Knowledge Distillation module aimed at distilling knowledge into the student network [38, 62].
To maximize the generalizability of the pretrained model, it is crucial to balance the performance and diversity of expert models. We evaluated several existing models across six different tasks, selecting those that excelled in classification (UNI), report generation (Phikon), and visual question answering (CONCH) as expert models (see Fig. 1c).
The [CLS] token, which represents the overall information of a patch for downstream tasks, serves as a critical component in our approach.
If the [CLS] token of our model aligns well with those of the expert models, it indicates that our model can effectively assimilate the knowledge from selected experts.
Similarly, the [PATCH] token also contains rich information.
For example, some methods use mean pooling to perform downstream tasks [63].
Therefore, aligning the [PATCH] token can further improve the effect of knowledge transfer.
To achieve above alignments, we use the student network to encode the global views and and extract the [CLS] and [PATCH] tokens.
Additionally, we employ the adopted experts to obtain their [CLS] and [PATCH] tokens, respectively.
For aligning the class tokens, we utilize cosine similarity. As for the patch token alignment, we employ both cosine similarity and smooth L1 distance. The pseudo-code for this process is outlined in Algorithm 1.
@@
Baselines. To evaluate the performance of our FM, GPFM, we conducted a comprehensive evaluation by comparing it with other vision FMs, namely Ctranspath [37], Phikon [32], and UNI [33], slide-level FM CHIEF [64] and Prov-Gigapath [56], as well as visual-language FMs PLIP [36] and CONCH [35].
As a baseline, we also compared these FMs with a ResNet50 [65] pretrained on the ImageNet dataset [66].
@@
For all downstream tasks, it should be emphasized that feature extraction was consistently performed on images resized to 224×224 resolution, except where explicitly stated otherwise in the experimental protocol.